# Weighted Birkhoff averages: Deterministic and probabilistic perspectives

### v2

@@ 3 Proof of the deterministic results
The result (I)–(1) was established in [DSSY17, Theorem 3.1], [DY18, Theorem 1.1], [DM23, Theorem 3], and [TL24a, Theorem 2.1]. The result (I)–(2) was initially obtained in [TL24a, Corollary 2.1] in a qualitative form, where the exponent $\zeta_{1}>0$ in $\mathcal{O}\left(\exp(-N^{\zeta_{1}})\right)$ is small and not explicitly determined. It was subsequently derived in [TL25a, Theorem 6.1], in a qualitative manner. For $p,q>0$, the result (I)–(2) can be derived as a byproduct from the proof of Theorem 1.4, see Section 3.3 for details. The result (I)–(3) was proved in [TL25a, Theorem 6.2]. The proof of the result (I)–(4) is similar to that of (III)–(7), which will be discussed later. For $p=q=1$, the result (II)–(5) was initially obtained in [TL24a, Corollary 3.1] in a multiple version and in a qualitative form. Similarly, for the quantitative result regarding $p,q>0$, it can also be derived from the proof of Theorem 1.4. A weaker version of the result (II)–(6) was presented in [TL24a, Corollary 3.2] and [TL24c, Theorem 4.4]. Given that the stronger conclusion can be derived by similar arguments (also combining the approach in Theorem 1.4), we omit the proof here for brevity. Although some of the aforementioned references only discuss the discrete case, the proof for the continuous case is relatively more straightforward, as pointed out in [TL24a, TL24c] and others. Consequently, we provide the proof only for the result (III)–(7) in this section.
Consider the $T$-periodic case with $T\in\mathbb{N}^{+}$, namely the data $f(0),f(1),\cdots,f(T-1)$ with $f$ being $2\pi$-periodic. We first need a quantitative version of the inverse discrete Fourier theorem. As it can be seen later, the order of the trigonometric polynomial adapted to the data is somewhat important, since we have to avoid the small divisor to achieve exponential convergence.
@@
$f\left(n\right)={a_{0}}+\sum\limits_{k=1}^{m}{{a_{k}}\cos\left({k\pi n/{m}}\right)}+\sum\limits_{k=1}^{m-1}{{b_{k}}\sin\left({k\pi n/{m}}\right)},\;\;1\leqslant n\leqslant 2m.$ (3.7)
@@
$f\left(n\right)={a_{0}}+\sum\limits_{k=1}^{m}{\left({{a_{k}}\cos\left({2k\pi n/{{(2m+1)}}}\right)+{b_{k}}\sin\left({2k\pi n/{{(m+1)}}}\right)}\right)},\;\;1\leqslant n\leqslant 2m+1.$ (3.8)
@@
Then the equations in (3.7) with all $1\leqslant n\leqslant 2m$ yield $f=\mathscr{V}\mathscr{C}$. We can observe the orthogonality of vectors in $\mathscr{V}$, i.e., $\mathscr{V}_{l}\cdot\mathscr{V}_{j}=0$ for all $l\neq j$. For example, define ${\alpha_{1}}=:\pi\left({l+j}\right)/m$ and ${\alpha_{2}}=:\pi\left({l-j}\right)/m$ for $1\leqslant l\leqslant m$ and $1\leqslant j\leqslant m-1$, then it holds
@@
$=\frac{1}{2}\left({\frac{{\sin\left({\left({2m-1}\right){\alpha_{1}}/2}\right)}}{{\sin\left({{\alpha_{1}}/2}\right)}}\sin\left({m{\alpha_{1}}}\right)-\frac{{\sin\left({\left({2m-1}\right){\alpha_{2}}/2}\right)}}{{\sin\left({{\alpha_{2}}/2}\right)}}\sin\left({m{\alpha_{2}}}\right)}\right)=0$
due to $\sin\left({m{\alpha_{1}}}\right)=\sin\left({m{\alpha_{2}}}\right)=0$ (we define $\sin\left({qx}\right)/\sin x=:q$ for $x=0$). For orthogonality of other column vectors in $\mathscr{V}$, see Appendix 5.1. Note that the orthogonality implies the uniqueness of $\mathscr{C}$, so the $2m$-periodic data can be expressed as a linear superposition of trigonometric polynomial of up to order $m$. Finally, by summing up all the row vectors in $\mathscr{V}$ and using
$\sum\limits_{n=0}^{2m-1}{{\exp\left(l\pi ni/m\right)}}=\frac{{1-{\exp\left(2l\pi i\right)}}}{{1-{\exp\left({l\pi i/m}\right)}}}=0,\;\;\forall 1\leqslant l\leqslant m,$ (3.9)
we obtain that ${a_{0}}={T^{-1}}\sum\nolimits_{n=0}^{T-1}{f\left(n\right)}$.
@@
$f\left(n\right)={a_{0}}+\sum\limits_{1\leqslant\left|k\right|\leqslant\mathscr{M}_{T}}{{c_{k}}{\exp\left(2k\pi ni/T\right)}},\;\;\forall n\in\mathbb{N},$ (3.10)
where $c_{k}$ are complex coefficients. Note that for all $1\leqslant\left|k\right|\leqslant\mathscr{M}\left(T\right)$, we have $\left|{k/T}\right|\leqslant\left|{\mathscr{M}_{T}/T}\right|\leqslant\left|{\left({T/2}\right)/T}\right|=1/2$. This leads to $\left|{k/T-n}\right|\geqslant 1/2$ for all $n\in\mathbb{Z}$. In other words, no small divisors will appear in the following analysis (for more details, see also [TL24c, Section 3.2]). Then we can conclude from the Poisson summation formula [Gra14, Chapter 3] that
@@
which admits exponential smallness $\mathcal{O}\left({\exp(-cN^{{\zeta_{5}}})}\right)$ for some universal constant $c>0$, where ${\zeta_{5}}={\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$, by using similar approach in [TL25a]. Alternatively, we can utilize the main ideas from the proof of Theorem 1.4 in Section 3.3. The distinction lies in the fact that the $m_{N}$ defined in (3.17) is replaced with ${m_{N}}\sim{N^{1/{\beta_{p,q}}}}$ as $N\to+\infty$, where ${\beta_{p,q}}:=1+1/\min\left\{{p,q}\right\}>1$, and the convergence rate in (3.18) is replaced with $\mathcal{O}\left({\exp(-cN^{{\zeta_{5}}})}\right)$. Finally, in view of (3.10), we obtain the exponential convergence for the weighted $T$-periodic data to its periodic average as
@@
provided implied constants that only depend on $T$ and ${\max_{0\leqslant n\leqslant T}}\left|{f\left(n\right)}\right|$. ∎
@@
$\hat{K}\left(s\right)$ $=\int_{{\mathbb{T}^{d}}}{K\left(x\right)\exp\left({-2\pi i\left\langle{s,x}\right\rangle}\right)dx}$
$=\mathop{\lim}\limits_{N\to+\infty}\frac{1}{N}\sum\limits_{n=1}^{N-1}{K\left({{\theta_{n}}}\right)\exp\left({-2\pi i\left\langle{s,{\theta_{n}}}\right\rangle}\right)}\;\;\;\;(\text{recall that ${\theta_{n}}:={\theta_{0}}+n\rho$})$
@@
$\sum\limits_{s\neq k\in{\mathbb{Z}^{d}}(\mathbb{Z}_{*}^{\infty})}{\hat{K}\left(k\right)\exp\left({2\pi i\left\langle{k-s,{\theta_{0}}}\right\rangle}\right)\left({\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right)\exp\left({2\pi in\left\langle{k-s,\rho}\right\rangle}\right)}}\right)}.$
@@
${\mathscr{E}_{N}}:=\sum\limits_{s\neq k\in{\mathbb{Z}^{d}}(\mathbb{Z}_{*}^{\infty})}{{{\|{\hat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}\left|{\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right)\exp\left({2\pi in\left\langle{k-s,\rho}\right\rangle}\right)}}\right|}.$ (3.13)
@@
$\left|{\left\langle{k,\rho}\right\rangle-n}\right|>\gamma\left\|k\right\|_{{\ell^{1}}}^{-\tau},\;\;\gamma>0,\;\;\forall 0\neq k\in\mathbb{Z}^{d},\;\;n\in\mathbb{Z},$
@@
${\mathcal{S}_{\rm T}^{d}}:=\left\{{k\in{\mathbb{Z}^{d}}:1\leqslant{{\left\|{k-s}\right\|}_{{\ell^{1}}}}\leqslant{N^{1/\left({\tau+{\beta_{p,q}}}\right)}}}\right\},$
@@
${\mathcal{S}_{\rm R}^{d}}:=\left\{{k\in{\mathbb{Z}^{d}}:{{\left\|{k-s}\right\|}_{{\ell^{1}}}}>{N^{1/\left({\tau+{\beta_{p,q}}}\right)}}}\right\}.$
@@
${\mathscr{E}_{N}}$ $=\sum\limits_{k\in{\mathcal{S}_{\rm T}^{d}}}{+\sum\limits_{k\in{\mathcal{S}_{\rm R}^{d}}}{{{\|{\hat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}\left|{\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right)\exp\left({2\pi in\left\langle{k-s,\rho}\right\rangle}\right)}}\right|}}$
@@
Based on the observation that for any fixed $k\in\mathcal{S}_{\rm R}^{d}$ and nonresonant $\rho\in\mathbb{R}^{d}$, there exists a unique $n^{*}\in\mathbb{Z}$ such that $\left|{\left\langle{k-s,\rho}\right\rangle-{n^{*}}}\right|<1/2$ which leads to small divisors, we obtain from [TL25a, Lemma 4.1] (the asymptotic estimate of $\left\|{{D^{m}}{w_{p,q}}}\right\|_{{L^{1}}\left({0,1}\right)}$} for $m$ sufficiently large) that
@@
${m_{N}}\sim\exp\left({-1}\right)\lambda_{*}^{-1/{\beta_{p,q}}}{N^{1/\left({\tau+{\beta_{p,q}}}\right)}}\to+\infty,\;\;m_{N}\in\mathbb{N}^{+},$ (3.17)
@@
$\lesssim\exp\left({-{{\beta_{p,q}}}{m_{N}}}\right)\lesssim{\exp(-c{N^{1/\left({\tau+{\beta_{p,q}}}\right)}})},$ (3.18)
@@
$\left|{\left\langle{k-s,\rho}\right\rangle-n}\right|\geqslant\left|{n-{n^{*}}}\right|-\left|{\left\langle{k-s,\rho}\right\rangle-{n^{*}}}\right|>\left|{n-{n^{*}}}\right|-1/2,\;\;\forall{n^{*}},{n^{*}}\pm 1\neq n\in\mathbb{Z}.$
Now, substituting (3.18) and (3.19) into (3.15) and utilizing the simple facts $\sum\nolimits_{k\in{\mathcal{S}_{\rm T}}}{{{\|{\hat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}}\leqslant\sum\nolimits_{k\in\mathbb{Z}^{d}}{{{\|{\hat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}}<+\infty$‡‡‡‡‡‡It is evident that the analyticity of $K$ implies the absolute summability of its Fourier coefficients. and $N/A_{N}^{p,q}\sim{({\int_{0}^{1}{{w_{p,q}}\left(x\right)dx}})^{-1}}=1$, we obtain
@@
${N^{1/\left({\tau+{\beta_{p,q}}}\right)}}<{\left\|{k-s}\right\|_{{\ell^{1}}}}\leqslant{\left\|k\right\|_{{\ell^{1}}}}+{\left\|s\right\|_{{\ell^{1}}}},\;\;k\in{\mathcal{S}_{\rm R}^{d}},$
@@
$\mathscr{E}_{N}^{(2)}$ $\leqslant\sum\limits_{k\in\mathcal{S}_{\rm R}^{d}}{{{\|{\hat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}}\leqslant{\left\|K\right\|_{\sigma,d}}\sum\limits_{k\in\mathcal{S}_{\rm R}^{d}}{\exp\left({-2\pi\sigma{{\left\|k\right\|}_{{\ell^{1}}}}}\right)}$
@@
It is observed that the truncation technique employed here can eliminate the nonresonance between $\rho$ and $k$ within the residual set $\mathcal{S}_{\rm R}^{d}$}, thereby avoiding the need for small divisor estimates similar to those in $\mathscr{E}_{N}^{(1)}$. This idea, initially appeared in [TL24c], prevents the increase in error, as small divisors can significantly undermine the controllability brought by analyticity as $\|k\|_{\ell^{1}}$ increases to $\infty$.
@@
$\mathscr{E}_{N}\lesssim{\exp(-{N^{{\zeta_{6}}}})},\;\;N\to+\infty$
@@
$\left|{\left\langle{k,\rho}\right\rangle-n}\right|>\gamma{\left({\prod\nolimits_{j\in{\mathbb{N}}}{(1+{{\left\langle j\right\rangle}^{\tau}}{{\left|{{k_{j}}}\right|}^{\tau}})}}\right)^{-1}},\;\;\gamma>0,\;\;\forall 0\neq k\in\mathbb{Z}^{\infty}_{*},\;\;n\in\mathbb{Z},$ (3.22)
@@
$\mathcal{S}_{\rm T}^{\infty}:=\left\{{k\in\mathbb{Z}_{*}^{\infty}:1\leqslant{{\left|{k-s}\right|}_{\eta}}\leqslant{{({\log N})}^{\zeta_{7}}}}\right\},$
@@
$\mathcal{S}_{\rm R}^{\infty}:=\left\{{k\in\mathbb{Z}_{*}^{\infty}:{{\left|{k-s}\right|}_{\eta}}>{{({\log N})}^{\zeta_{7}}}}\right\}.$
@@
${{\tilde{m}}_{N}}\sim\exp\left({-1}\right){N^{1/{\beta_{p,q}}}}\exp\left({\beta_{p,q}^{-1}c{{(\log N)}^{\frac{{\zeta_{7}}}{{1+\eta}}}}\log\log N}\right),\;\;m_{N}\in\mathbb{N}^{+},$
@@
${{\tilde{m}}_{N}}\gtrsim{(\log N)^{\zeta_{7}}},\;\;{N^{a}}\gtrsim{(\log N)^{\zeta_{7}}}\;\text{for any $a>0$}$
@@
${\tilde{\mathcal{E}}_{N}^{(2)}}$ $\lesssim\sum\limits_{k\in\mathcal{S}_{\rm R}^{\infty}}{{{\|{\hat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}}\leqslant{\left\|K\right\|_{\sigma,\infty}}\sum\limits_{k\in\mathcal{S}_{\rm R}^{\infty}}{\exp\left({-2\pi\sigma{{\left|k\right|}_{\eta}}}\right)}$
@@
Next, we need to prove the previous claim (3.28). In this case, distinct from the finite-dimensional setting, the infinite-dimensional spatial structure compels us to derive cardinality estimates on the infinite-dimensional lattice, as initially observed in [TL24a, TL24c]. Recall that [TL24c, Lemma 8.4] gives
$\sum\limits_{k\in\mathbb{Z}_{*}^{\infty},{{\left|k\right|}_{\eta}}=\nu}1:=\#\left\{{k:\;\;k\in\mathbb{Z}_{*}^{\infty},\;{{\left|k\right|}_{\eta}}=\nu\in{\mathbb{N}^{+}}}\right\}\lesssim{\nu^{{\nu^{1/\eta}}}},\;\;\nu\to+\infty.$
@@
${\mathscr{E}_{N}}\lesssim{\tilde{\mathcal{E}}_{N}^{(1)}}+{\tilde{\mathcal{E}}_{N}^{(2)}}\lesssim{\exp({-{{(\log N)}^{\zeta_{7}}}})},\;\;N\to+\infty$
@@
As a remark, the universal infinite-dimensional nonresonance beyond (3.22) was also constructed in [TL24c], and the exponential convergence was proved via analyticity (i.e., two distinct approaches were presented there). The method in this section can also be employed to enhance the other approach in [TL24c].

### v3

@@ 3 Proof of the deterministic results
The result (I)–(1) was established in [DSSY17, Theorem 3.1], [DY18, Theorem 1.1], [DM23, Theorem 3], and [TL24a, Theorem 2.1]. The result (I)–(2) was initially obtained in [TL24a, Corollary 2.1] in a qualitative form, where the exponent $\zeta_{1}>0$ in $\mathcal{O}\left(\exp(-N^{\zeta_{1}})\right)$ is small and not explicitly determined. It was subsequently derived in [TL25b, Theorem 6.1], in a qualitative manner. For $p,q>0$, the result (I)–(2) can be derived as a byproduct from the proof of Theorem 1.4, see Section 3.3 for details. The result (I)–(3) was proved in [TL25b, Theorem 6.2]. The proof of the result (I)–(4) is similar to that of (III)–(7), which will be discussed later. For $p=q=1$, the result (II)–(5) was initially obtained in [TL24a, Corollary 3.1] in a multiple version and in a qualitative form. Similarly, for the quantitative result regarding $p,q>0$, it can also be derived from the proof of Theorem 1.4. A weaker version of the result (II)–(6) was presented in [TL24a, Corollary 3.2] and [TL24b, Theorem 4.4]. Given that the stronger conclusion can be derived by similar arguments (also combining the approach in Theorem 1.4), we omit the proof here for brevity. Although some of the aforementioned references only discuss the discrete case, the proof for the continuous case is relatively more straightforward, as pointed out in [TL24a, TL24b] and others. Consequently, we provide the proof only for the result (III)–(7) in this section.
Consider the $T$-periodic case with $T\in\mathbb{N}^{+}$, namely the data $f(0),f(1),\cdots,f(T-1)$ with $f$ being $T$-periodic. We first need a quantitative version of the inverse discrete Fourier theorem. As it can be seen later, the order of the trigonometric polynomial adapted to the data is somewhat important, since we have to avoid the small divisor to achieve exponential convergence.
@@
$f\left(n\right)={a_{0}}+\sum\limits_{k=1}^{m}{{a_{k}}\cos\left({k\pi n/{m}}\right)}+\sum\limits_{k=1}^{m-1}{{b_{k}}\sin\left({k\pi n/{m}}\right)},\quad 1\leqslant n\leqslant 2m.$ (3.7)
@@
$f\left(n\right)={a_{0}}+\sum\limits_{k=1}^{m}{\left({{a_{k}}\cos\left({2k\pi n/{{(2m+1)}}}\right)+{b_{k}}\sin\left({2k\pi n/{{(m+1)}}}\right)}\right)},\quad 1\leqslant n\leqslant 2m+1.$ (3.8)
@@
Then the equations in (3.7) with all $1\leqslant n\leqslant 2m$ yield $f=\mathscr{V}\mathscr{C}$.
We now establish the orthogonality of vectors in $\mathscr{V}$, namely $\mathscr{V}_{l}\cdot\mathscr{V}_{j}=0$ whenever $l\neq j$. Set ${\alpha_{1}}:=\pi\left({l+j}\right)/m$ and ${\alpha_{2}}:=\pi\left({l-j}\right)/m$ for $1\leqslant l\leqslant m$ and $1\leqslant j\leqslant m-1$; then
@@
$=\frac{1}{2}\left({\frac{{\sin\left({\left({2m-1}\right){\alpha_{1}}/2}\right)}}{{\sin\left({{\alpha_{1}}/2}\right)}}\sin\left({m{\alpha_{1}}}\right)-\frac{{\sin\left({\left({2m-1}\right){\alpha_{2}}/2}\right)}}{{\sin\left({{\alpha_{2}}/2}\right)}}\sin\left({m{\alpha_{2}}}\right)}\right)=0,$
because $\sin\left({m{\alpha_{1}}}\right)=\sin\left({m{\alpha_{2}}}\right)=0$, with the convention $\sin\left({qx}\right)/\sin x:=q$ when $x=0$. For $\mathscr{V}_{2l-1}$} and $\mathscr{V}_{2j-1}$} with $l\neq j$ and $l+j\leqslant m+\left({m-1}\right)=2m-1<2m$, we obtain from ${\alpha_{1}},{\alpha_{2}}\in\left({0,\pi}\right)$ that
An analogous computation gives ${\mathscr{V}_{2l}}\cdot{\mathscr{V}_{2j}}=0$ for $l\neq j$ and $l+j\leqslant 2m-2<2m$:
Finally, the identity
$\sum\limits_{n=0}^{2m-1}{{\exp\left(l\pi ni/m\right)}}=\frac{{1-{\exp\left(2l\pi i\right)}}}{{1-{\exp\left({l\pi i/m}\right)}}}=0,\quad\forall 1\leqslant l\leqslant m$ (3.9)
provides the orthogonality of $\mathscr{V}_{0}$ and other vectors.
Note that the orthogonality implies the uniqueness of $\mathscr{C}$, so the $2m$-periodic data can be expressed as a linear superposition of trigonometric polynomial of up to order $m$. Finally, by summing up all the row vectors in $\mathscr{V}$ and using (3.9), we obtain that ${a_{0}}={T^{-1}}\sum\nolimits_{n=0}^{T-1}{f\left(n\right)}$.
@@
$f\left(n\right)={a_{0}}+\sum\limits_{1\leqslant\left|k\right|\leqslant\mathscr{M}_{T}}{{c_{k}}{\exp\left(2k\pi ni/T\right)}},\quad\forall n\in\mathbb{N},$ (3.10)
where $c_{k}$ are complex coefficients. Note that for all $1\leqslant\left|k\right|\leqslant\mathscr{M}\left(T\right)$, we have $\left|{k/T}\right|\leqslant\left|{\mathscr{M}_{T}/T}\right|\leqslant\left|{\left({T/2}\right)/T}\right|=1/2$. This leads to $\left|{k/T-n}\right|\geqslant 1/2$ for all $n\in\mathbb{Z}$. In other words, no small divisors will appear in the following analysis (for more details, see also [TL24b, Section 3.2]). Then we can conclude from the Poisson summation formula [Gra14, Chapter 3] that
@@
which admits exponential smallness $\mathcal{O}\left({\exp(-cN^{{\zeta_{5}}})}\right)$ for some universal constant $c>0$, where ${\zeta_{5}}={\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$, by using similar approach in [TL25b]. Alternatively, we can utilize the main ideas from the proof of Theorem 1.4 in Section 3.3. The distinction lies in the fact that the $m_{N}$ defined in (3.17) is replaced with ${m_{N}}\sim{N^{1/{\beta_{p,q}}}}$ as $N\to+\infty$, where ${\beta_{p,q}}:=1+1/\min\left\{{p,q}\right\}>1$, and the convergence rate in (3.18) is replaced with $\mathcal{O}\left({\exp(-cN^{{\zeta_{5}}})}\right)$. Finally, in view of (3.10), we obtain the exponential convergence for the weighted $T$-periodic data to its periodic average as
@@
provided implied constants that only depend on $T$ and ${\max_{0\leqslant n\leqslant T}}\left|{f\left(n\right)}\right|$ (here we regard $p$ and $q$ as fixed). ∎
@@
$\widehat{K}\left(s\right)$ $=\int_{{\mathbb{T}^{d}}}{K\left(x\right)\exp\left({-2\pi i\left\langle{s,x}\right\rangle}\right)dx}$
$=\mathop{\lim}\limits_{N\to+\infty}\frac{1}{N}\sum\limits_{n=1}^{N-1}{K\left({{\theta_{n}}}\right)\exp\left({-2\pi i\left\langle{s,{\theta_{n}}}\right\rangle}\right)}\quad(\text{recall that ${\theta_{n}}:={\theta_{0}}+n\rho$})$
@@
$\sum\limits_{s\neq k\in{\mathbb{Z}^{d}}(\mathbb{Z}_{*}^{\infty})}{\widehat{K}\left(k\right)\exp\left({2\pi i\left\langle{k-s,{\theta_{0}}}\right\rangle}\right)\left({\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right)\exp\left({2\pi in\left\langle{k-s,\rho}\right\rangle}\right)}}\right)}.$
@@
${\mathscr{E}_{N}}:=\sum\limits_{s\neq k\in{\mathbb{Z}^{d}}(\mathbb{Z}_{*}^{\infty})}{{{\|{\widehat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}\left|{\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right)\exp\left({2\pi in\left\langle{k-s,\rho}\right\rangle}\right)}}\right|}.$ (3.13)
@@
$\left|{\left\langle{k,\rho}\right\rangle-n}\right|>\gamma\left\|k\right\|_{{\ell^{1}}}^{-\tau},\quad\gamma>0,\quad\forall 0\neq k\in\mathbb{Z}^{d},\quad\forall n\in\mathbb{Z},$
@@
${\mathcal{S}_{\rm T}^{d}}:=\left\{{k\in{\mathbb{Z}^{d}}:\quad 1\leqslant{{\left\|{k-s}\right\|}_{{\ell^{1}}}}\leqslant{N^{1/\left({\tau+{\beta_{p,q}}}\right)}}}\right\},$
@@
${\mathcal{S}_{\rm R}^{d}}:=\left\{{k\in{\mathbb{Z}^{d}}:\quad{{\left\|{k-s}\right\|}_{{\ell^{1}}}}>{N^{1/\left({\tau+{\beta_{p,q}}}\right)}}}\right\}.$
@@
${\mathscr{E}_{N}}$ $=\sum\limits_{k\in{\mathcal{S}_{\rm T}^{d}}}{+\sum\limits_{k\in{\mathcal{S}_{\rm R}^{d}}}{{{\|{\widehat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}\left|{\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right)\exp\left({2\pi in\left\langle{k-s,\rho}\right\rangle}\right)}}\right|}}$
@@
Based on the observation that for any fixed $k\in\mathcal{S}_{\rm R}^{d}$ and nonresonant $\rho\in\mathbb{R}^{d}$, there exists a unique $n^{*}\in\mathbb{Z}$ such that $\left|{\left\langle{k-s,\rho}\right\rangle-{n^{*}}}\right|<1/2$ which leads to small divisors, we obtain from [TL25b, Lemma 4.1] (the asymptotic estimate of $\left\|{{D^{m}}{w_{p,q}}}\right\|_{{L^{1}}\left({0,1}\right)}$} for $m$ sufficiently large) that
@@
${m_{N}}\sim\exp\left({-1}\right)\lambda_{*}^{-1/{\beta_{p,q}}}{N^{1/\left({\tau+{\beta_{p,q}}}\right)}}\to+\infty,\quad m_{N}\in\mathbb{N}^{+},$ (3.17)
@@
$\lesssim\exp\left({-{{\beta_{p,q}}}{m_{N}}}\right)\lesssim{\exp\left(-c{N^{1/\left({\tau+{\beta_{p,q}}}\right)}}\right)},$ (3.18)
@@
$\left|{\left\langle{k-s,\rho}\right\rangle-n}\right|\geqslant\left|{n-{n^{*}}}\right|-\left|{\left\langle{k-s,\rho}\right\rangle-{n^{*}}}\right|>\left|{n-{n^{*}}}\right|-1/2,\quad\forall{n^{*}},{n^{*}}\pm 1\neq n\in\mathbb{Z}.$
Now, substituting (3.18) and (3.19) into (3.15) and utilizing the simple facts $\sum\nolimits_{k\in{\mathcal{S}_{\rm T}}}{{{\|{\widehat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}}\leqslant\sum\nolimits_{k\in\mathbb{Z}^{d}}{{{\|{\widehat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}}<+\inftyIt$ is evident that the analyticity of $K$ implies the absolute summability of its Fourier coefficients. and $N/A_{N}^{p,q}\sim{({\int_{0}^{1}{{w_{p,q}}\left(x\right)dx}})^{-1}}=1$, we obtain
@@
${N^{1/\left({\tau+{\beta_{p,q}}}\right)}}<{\left\|{k-s}\right\|_{{\ell^{1}}}}\leqslant{\left\|k\right\|_{{\ell^{1}}}}+{\left\|s\right\|_{{\ell^{1}}}},\quad k\in{\mathcal{S}_{\rm R}^{d}},$
@@
$\mathscr{E}_{N}^{(2)}$ $\leqslant\sum\limits_{k\in\mathcal{S}_{\rm R}^{d}}{{{\|{\widehat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}}\leqslant{\left\|K\right\|_{\sigma,d}}\sum\limits_{k\in\mathcal{S}_{\rm R}^{d}}{\exp\left({-2\pi\sigma{{\left\|k\right\|}_{{\ell^{1}}}}}\right)}$
@@
It is observed that the truncation technique employed here can eliminate the nonresonance between $\rho$ and $k$ within the residual set $\mathcal{S}_{\rm R}^{d}$}, thereby avoiding the need for small divisor estimates similar to those in $\mathscr{E}_{N}^{(1)}$. This idea, initially appeared in [TL24b], prevents the increase in error, as small divisors can significantly undermine the controllability brought by analyticity as $\|k\|_{\ell^{1}}$ increases to $\infty$.
@@
$\mathscr{E}_{N}\lesssim{\exp(-{N^{{\zeta_{6}}}})},\quad N\to+\infty$
@@
$\left|{\left\langle{k,\rho}\right\rangle-n}\right|>\gamma{\left({\prod\nolimits_{j\in{\mathbb{N}}}{(1+{{\left\langle j\right\rangle}^{\tau}}{{\left|{{k_{j}}}\right|}^{\tau}})}}\right)^{-1}},\quad\gamma>0,\quad\forall 0\neq k\in\mathbb{Z}^{\infty}_{*},\quad\forall n\in\mathbb{Z},$ (3.22)
@@
$\mathcal{S}_{\rm T}^{\infty}:=\left\{{k\in\mathbb{Z}_{*}^{\infty}:\quad 1\leqslant{{\left|{k-s}\right|}_{\eta}}\leqslant{{({\log N})}^{\zeta_{7}}}}\right\},$
@@
$\mathcal{S}_{\rm R}^{\infty}:=\left\{{k\in\mathbb{Z}_{*}^{\infty}:\quad{{\left|{k-s}\right|}_{\eta}}>{{({\log N})}^{\zeta_{7}}}}\right\}.$
@@
${{\tilde{m}}_{N}}\sim\exp\left({-1}\right){N^{1/{\beta_{p,q}}}}\exp\left({\beta_{p,q}^{-1}c{{(\log N)}^{\frac{{\zeta_{7}}}{{1+\eta}}}}\log\log N}\right),\quad{\tilde{m}}_{N}\in\mathbb{N}^{+},$
@@
${{\tilde{m}}_{N}}\gtrsim{(\log N)^{\zeta_{7}}},\quad{N^{a}}\gtrsim{(\log N)^{\zeta_{7}}}\quad\text{for any $a>0$}$
@@
${\tilde{\mathcal{E}}_{N}^{(2)}}$ $\lesssim\sum\limits_{k\in\mathcal{S}_{\rm R}^{\infty}}{{{\|{\widehat{K}\left(k\right)}\|}_{{\ell^{\infty}}}}}\leqslant{\left\|K\right\|_{\sigma,\infty}}\sum\limits_{k\in\mathcal{S}_{\rm R}^{\infty}}{\exp\left({-2\pi\sigma{{\left|k\right|}_{\eta}}}\right)}$
@@
Next, we need to prove the previous claim (3.28). In this case, distinct from the finite-dimensional setting, the infinite-dimensional spatial structure compels us to derive cardinality estimates on the infinite-dimensional lattice, as initially observed in [TL24a, TL24b]. Recall that [TL24b, Lemma 8.4] gives
$\sum\limits_{k\in\mathbb{Z}_{*}^{\infty},{{\left|k\right|}_{\eta}}=\nu}1:=\#\left\{{k:\quad k\in\mathbb{Z}_{*}^{\infty},\;{{\left|k\right|}_{\eta}}=\nu\in{\mathbb{N}^{+}}}\right\}\lesssim{\nu^{{\nu^{1/\eta}}}},\quad\nu\to+\infty.$
@@
${\mathscr{E}_{N}}\lesssim{\tilde{\mathcal{E}}_{N}^{(1)}}+{\tilde{\mathcal{E}}_{N}^{(2)}}\lesssim{\exp({-{{(\log N)}^{\zeta_{7}}}})},\quad N\to+\infty$
@@
As a remark, the universal infinite-dimensional nonresonance beyond (3.22) was also constructed in [TL24b], and the exponential convergence was proved via analyticity (i.e., two distinct approaches were presented there). The method in this section can also be employed to enhance the other approach in [TL24b].